OpenStudy (anonymous):

ne~ can somebody please help me in solving projectile motion-related problems? i can't seem to identify whether the given is velocity (initial or final) in x or in y. thank you~ ^^

OpenStudy (anonymous):

it would be easier to explain if u post a problem

OpenStudy (anonymous):

in horizontal projectile motion the initial velocity is in x only in y direction the object is considered as a freely falling body in oblique projectile motion u(in x)=\[u \cos \theta\] \[u(\in y)=u \sin \theta\]

OpenStudy (anonymous):

hope u 'll understand